OpenStudy (shubhamsrg):

what is the maximum value for \[4/(4-x^2) + 9/(9-y^2)\] given that both \[x,y \in (-2,2)\] and xy=-1

OpenStudy (anonymous):

Use lagrange multipliers

OpenStudy (anonymous):

U know how to diffrentiate right

OpenStudy (shubhamsrg):

thats so very cool! thanx buddy..never knew about this.. and yes,,i know how to differentiaite..

OpenStudy (shubhamsrg):

differentiate*

OpenStudy (shubhamsrg):

leme try to solve here..

OpenStudy (shubhamsrg):

g(x,y) = xy+1 f(x,y) = 4/(4-x^2) + 9/(9-y^2) q(x,y,k) = f(x,y) + k(g(x,y)) = 4/(4-x^2) + 9/(9-y^2) + kxy + k dq/dx = 4(2x)/(4-x^2)^2 + ky dq/dy = 9(2y)/(9-y^2)^2 + kx dq/dk = xy + 1 on equating each to 0 , dont you think it'll be quite tedius to solve for x,y and k from here?

OpenStudy (shubhamsrg):

it gives quartic eqn in x and y.. it'll require to first convert it into cubic eqn by subsn,,then solve that eqn by converting to depressed eqn and make more substn..sir,,this is just a 4 mark objective ques..there has to be a simpler way..

OpenStudy (anonymous):

Dude g(x)=xy+1 del(f)=k.del(g) g=1 will give 3 equationa and 3 unknowns

OpenStudy (shubhamsrg):

didnt get you..

OpenStudy (anonymous):

\[\frac{df}{dx}i+\frac{df}{dy}j=\lambda[\frac{dg}{dx}i+\frac{dg}{dy}j]\] and \[xy=-1\]

OpenStudy (anonymous):

i solve it and got the answer

OpenStudy (shubhamsrg):

whats the ans?

OpenStudy (anonymous):

x=sqrt[29/21] and -sqrt[29/21]

OpenStudy (anonymous):

find y accodingly